OpenStudy (anonymous):

Choose the correct x-intercept & y-intercept of the equation 2x-y=4 ~ x-intercept : (-4,0) , y-intercept: (0,2) ~ x-intercept : (-4,0) , y-intercept: (2,0) ~ x-intercept : (2,0) , y-intercept: (0,-4) ~ x-intercept : (0,2) , y-intercept: (0,-4)

OpenStudy (anonymous):

Solve for y: 2 x = 4+y Reverse the equality in 2 x = 4+y in order to isolate y to the left hand side. 2 x = y+4 is equivalent to y+4 = 2 x: 4+y = 2 x Solve for y. Subtract 4 from both sides: Answer: | | y = -4+2 x

OpenStudy (anonymous):

Still dont understand .. I got 6

OpenStudy (anonymous):

@Luigi0210

OpenStudy (luigi0210):

x-intercept is when y=0 and y-intercept is when x=0

OpenStudy (luigi0210):

No giving out answers >.<

OpenStudy (anonymous):

srry

OpenStudy (anonymous):

So how do I figure out which points go with it without graphing it .

OpenStudy (anonymous):

standar equation of a line with x intercept (a,0) and y-intercept (0, b) xa+yb=1 here the equation is x+2y=4 divide the whole equation by 4 x4+y2=1

OpenStudy (luigi0210):

Plug in 0 for x to find the y: